OpenStudy (anonymous):

Determine the equation of the line, in standard form, that will get your spacecraft from the Launch Area to Point A. Launch area=(1, 2) Point A=(0, 3)

OpenStudy (anonymous):

It is on a graph from 0 in the middle and up to 5 on each of the four ends

OpenStudy (haseeb96):

first get the slope from both points okay do u know how to take out the slope from them

OpenStudy (anonymous):

m=y2-y1/x2-x1

OpenStudy (haseeb96):

yes apply it here

OpenStudy (anonymous):

m=3-2/0-1

OpenStudy (anonymous):

m=1/-1

OpenStudy (haseeb96):

yes m=-1 okay now use point slope form

OpenStudy (anonymous):

Ok, hold on

OpenStudy (anonymous):

Is it y+2=-1(x-1)

OpenStudy (anonymous):

y=-1x+0

OpenStudy (xapproachesinfinity):

your math is incorrect double check it

OpenStudy (xapproachesinfinity):

@LydiannAquino

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I'll be back hold on please.

OpenStudy (xapproachesinfinity):

okay

OpenStudy (anonymous):

y-(-2)=-1(x-1)

OpenStudy (xapproachesinfinity):

ok take off the parentheses

OpenStudy (anonymous):

y+2=-1x+2

OpenStudy (xapproachesinfinity):

y+2=-x+1 not -x+2

OpenStudy (xapproachesinfinity):

that was the mistake you made

OpenStudy (anonymous):

Oh, I see it

OpenStudy (xapproachesinfinity):

ok now we need standard form

OpenStudy (haseeb96):

y-y1=m(x-x1) y-2 =-1(x-1) y-2=-1x +1 y=-1x+1+2 y=-x +3 now in the standard from x+y-3=0 x must be +ve for standard form

OpenStudy (xapproachesinfinity):

oh he is right it should be y-2=-x+1 not what you wrote

OpenStudy (xapproachesinfinity):

so y=-x+3

OpenStudy (anonymous):

Ok.

OpenStudy (xapproachesinfinity):

x+y-3=0 , he already gave you standard form

OpenStudy (xapproachesinfinity):

ax+by+c=0

OpenStudy (anonymous):

The answer is y=-x+3 right

OpenStudy (anonymous):

and in slope intercept form it is y-2 =-1(x-1)

OpenStudy (xapproachesinfinity):

That is correct! however they asked for standard form y=-x+3 is slope intercept form and y-2=-(x-1) is point slope from not slope intercept form

OpenStudy (xapproachesinfinity):

the final correct answer is X+Y-3=0

OpenStudy (anonymous):

Is'nt standard form y=mx+b

OpenStudy (xapproachesinfinity):

No that's slope intercept form

OpenStudy (anonymous):

y=-x+3 is standard form nad slope intercept form is y-y1=m(x-x1)

OpenStudy (xapproachesinfinity):

a line is Ax+By+c in standard form

OpenStudy (xapproachesinfinity):

check here http://www.mathsisfun.com/algebra/linear-equations.html

OpenStudy (anonymous):

Ok so X+Y-3=0 that's standard form

OpenStudy (xapproachesinfinity):

Yep!

OpenStudy (anonymous):

Now, are you able to place it so I can see how we worked it all out?

OpenStudy (xapproachesinfinity):

Ax+By=c or Ax+By+c=0 you could see this as the standard form

OpenStudy (anonymous):

Yes, that's what I saw through the link you sent me.

OpenStudy (xapproachesinfinity):

alright let me write the summary of what we worked here

OpenStudy (anonymous):

Ok.

OpenStudy (xapproachesinfinity):

First we found the slope using the two given points (1,2) and (0,3) m=3-2/0-1=-1 Second we use the formula (y-y1)=m(x-x1) which is point slope form it doesn't matter what point you choose the answer will the same y-2=-1(x-1) ===> y-2=-x+1 ==> y=-x+3 we found the slope intercept form third step is the put x and y in one side to get the standard form that we want x+y=3

OpenStudy (xapproachesinfinity):

this is what we did!

OpenStudy (xapproachesinfinity):

I prefer to leave this way x+y=3 since textbooks do so

OpenStudy (anonymous):

so y=mx+b is not standard form I asked my teacher and she said it was correct but then... I'm really confused...??

OpenStudy (xapproachesinfinity):

No, it is not! that's slope intercept form. check your textbook

OpenStudy (anonymous):

Oh, no you're correct I just looked it up myself...sorry

OpenStudy (xapproachesinfinity):

No problem!

OpenStudy (anonymous):

y=-x+3 is slope intercept form, right? How did we get that answer for the next question?

OpenStudy (xapproachesinfinity):

Yes! we got it from this (y-y1)=m(x-x1)

OpenStudy (anonymous):

y-2=-(x-1) <this?

OpenStudy (xapproachesinfinity):

Yes!

OpenStudy (xapproachesinfinity):

we just took out the parentheses and subtracted 2 from both sides

OpenStudy (xapproachesinfinity):

I mean added 2 to both sides

OpenStudy (anonymous):

Ok I got it. Thank you very much, for everything

OpenStudy (xapproachesinfinity):

Pleasure to help you! Cheers

OpenStudy (anonymous):

Would any restrictions apply to the domain and range of this equation? Explain your reasoning using complete sentences. This is my last question could you help me?

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

I will be back,

OpenStudy (anonymous):

Ok.

OpenStudy (xapproachesinfinity):

So, is the question related to this function f(x)=-x+3 or something else?